# REAP the Experts: Why Pruning Prevails for One-Shot MoE compression

### v1

@@ Abstract
Sparsely-activated Mixture-of-Experts (SMoE) models offer efficient pre-training and low latency but their large parameter counts create significant memory overhead, motivating research into expert compression. Contrary to recent findings favouring expert merging on discriminative benchmarks, we demonstrate that expert pruning is a superior strategy for generative tasks. We prove that merging introduces an irreducible error by causing a “functional subspace collapse”, due to the loss of the router’s independent, input-dependent control over experts. Leveraging this insight, we propose Router-weighted Expert Activation Pruning (REAP), a novel pruning criterion that considers both router gate-values and expert activation norms. Across a diverse set of SMoE models ranging from 20B to 1T parameters, REAP consistently outperforms merging and other pruning methods on generative benchmarks, especially at 50% compression. Notably, our method achieves near-lossless compression on code generation and tool-calling tasks with Qwen3-Coder-480B and Kimi-K2, even after pruning 50% of experts.
@@ 1 Introduction
Interest in the Sparsely-activated Mixture-of-Experts (SMoE) architecture for Large Language Models (LLMs) surged following the release of DeepSeek-V3 (DeepSeek-AI et al., 2024) and other high-quality open-weight SMoE LLMs (Jiang et al., 2024; Meta AI Team, 2025; Yang et al., 2025a; Zeng et al., 2025; Baidu, 2025; Kimi Team et al., 2025). Compared to dense models, the SMoEs offer lower latency and more efficient pre-training (Fedus et al., 2022). However, SMoEs require more parameters than dense models to achieve similar accuracy, resulting in significant memory overhead. Further, expert usage imbalance during inference causes poor accelerator utilization, leading to increased latency or compromises such as dropped tokens (Balmau et al., 2025). Expert usage imbalance also represents an opportunity, motivating prior work which investigates whether experts can be compressed without negatively impairing accuracy (Li et al., 2023; Lu et al., 2024). By eliminating or compressing redundant experts, memory overhead is reduced. A more uniform distribution of expert usage would also improve hardware utilization. Expert compression is particularly valuable for use cases which feature small batch sizes such as local deployments and academic research.
@@
  • We prove that expert merging introduces irreducible error due to the loss of the router’s independent, input-dependant modulation of the expert outputs resulting in functional subspace collapse, substantially reducing the functional output space of the compressed SMoE layer. In contrast, in expert pruned SMoEs the router maintains independent control over the remaining experts;
  • We introduce Router-weighted Expert Activation Pruning (REAP), a novel expert pruning saliency criterion, which selects experts to prune which contribute minimally to the layer output by considering both the router gate-values and average activation norm of the experts;
  • Across diverse SMoE architectures ranging from 20B to 1T parameters and a suite of generative evaluations, we demonstrate the significant and consistent advantage of REAP over existing expert pruning and merging approaches, particularly at 50% compression. Notably, our method achieves near-lossless compression on code generation tasks after pruning 50% of experts from Qwen3-Coder-480B and Kimi-K2;
@@ 2 Related Work
A Mixture-of-Experts (MoE) layer is comprised of multiple, specialized feed-forward subnetworks known as experts and a router which produces gate-values (i.e., gates) to dynamically modulate the output of the experts based on the input. The architecture was revived in the deep learning era by the introduction of the SMoE by Shazeer et al. (2017). SMoEs layers only select a subset of experts to use for each input, enabling massive scaling of model parameters without a commensurate increase in computational cost (Lepikhin et al., 2021; Fedus et al., 2022). In transformer-based LLMs, SMoE layers are integrated by replacing the traditional feed-forward layers. Further innovations such as auxiliary-loss-free load balancing (DeepSeek-AI et al., 2024), shared experts, and fined-grained experts (Dai et al., 2024) have propelled SMoE architectures to become the de facto standard for LLMs in recent months.
@@
Although SMoE layers effectively decouple total model parameters from inference costs, the memory overhead of storing large SMoEs restricts their deployment in resourced-constrained environments, motivating research in expert pruning to reduce total number of parameters. Early efforts demonstrated that progressively pruning experts based on router weights during fine-tuning until a single expert remained could preserve model quality in task-specific settings (Chen et al., 2022). Koishekenov et al. (2023) found expert pruning to be effective without further fine-tuning despite aggressively pruning up to 80% of experts.  Muzio et al. (2024) found that global pruning using gate-values as a saliency criterion was more effective than uniform, layer-wise frequency-based pruning. Other sophisticated pruning criteria have been proposed:  Lu et al. (2024) introduced an exhaustive search strategy which prunes experts that minimize the reconstruction loss between the original and pruned layer outputs;  Liu et al. (2024a) used a gradient-free evolutionary algorithm to prune experts. Both of these works demonstrated significant improvements over naive frequency-based pruning. A comprehensive evaluation of 16 diverse pruning criteria was conducted by Jaiswal et al. (2025). Expert Activation Norm (EAN) was empirically found to be the highest performing criterion and the benefits of iterative pruning were presented.
@@
While the above-noted works prove that expert compression is feasible via pruning, an alternative compression technique is to merge experts. Generally, merging requires both a clustering algorithm and a merging technique. Li et al. (2023) introduced Merge Sparse Mixture of Experts (M-SMoE) which first initializes expert cluster centres by identifying the dominant experts with the highest usage frequency globally across all layers. The remaining non-dominant experts are clustered based on the cosine similarity of router logits. Finally, experts weights are aligned via permutation with the weight matching algorithm (Ainsworth et al., 2023) and merged using frequency-weighted parameter averaging.  Li et al. (2023) found that their technique outperformed Chen et al.’s (2022) pruning method on MC benchmarks. Chen et al. (2025) proposed Hierarchical Clustering for Sparsely activated Mixture of Experts (HC-SMoE). HC-SMoE clusters experts based on the euclidean similarity of their representative vectors — the average activation of each expert measured on every token in a calibration dataset — using hierarchical agglomerative clustering. Similar to M-SMoE, HC-SMoE uses frequency-weighted parameter averaging to merge clusters into a single merged expert. Without any fine-tuning, Chen et al. (2025) found that their technique outperformed expert pruning based on router logits (He et al., 2025a), frequency, and Lu et al.’s (2024) method when benchmarked on a suite of MC question answering tasks.
@@
In addition to pruning and merging, experts may be compressed through quantization (Huang et al., 2025), low-rank decomposition (Yang et al., 2024a; Gu et al., 2025; He et al., 2025b), weight sparsity (He et al., 2025a), or a combination of any of the above techniques (Liu et al., 2025). These other approaches are orthogonal to expert pruning and merging; however, note that expert merging necessitates re-quantization for block quantization formats that share common scaling coefficients across a group of weights.
@@
3 Merging Experts Causes Functional Subspace Collapse
@@
To motivate our proposed expert pruning method, we first formally develop the expected errors of both expert merging and pruning. Consider a SMoE layer with $K$ experts $f_{1},\dots,f_{K}$, each a function $f_{k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and a router producing non-negative gates $\mathbf{g(x)}=(g_{1}(x),\dots,g_{K}(x))\in\Delta^{K-1}$. Top-$k$ routing is achieved by zeroing all but the largest $k$ gates. The output of the original layer is
$h(x):=\sum_{k=1}^{K}g_{k}(x)\,f_{k}(x).$ (1)
@@
To analyse the fundamental difference between compression operations, we focus on the elementary case of reducing two experts, $(f_{i},f_{j})$, to one. This pairwise analysis is the building block for any larger merge within a cluster. Pruning removes expert $j$ and re-normalizes the router outputs over the remaining K1K-1 experts, producing a new set of gates $\bar{g}(x)$. Merging replaces $(f_{i},f_{j})$ with a new expert $\tilde{f}$. Existing one-shot expert merging methods such as HC-SMoE and M-SMoE sum the gates for the original experts $g_{i}(x)+g_{j}(x)$. The pruned, $\bar{h}(x)$, and merged, $\tilde{h}(x)$, layer outputs are
$\bar{h}(x):=\sum_{k\neq j}\bar{g}_{k}(x)\,f_{k}(x),$ (2)
$\tilde{h}(x):=\sum_{k\neq i,j}g_{k}(x)f_{k}(x)+\big(g_{i}(x)+g_{j}(x)\big)\tilde{f}(x).$ (3)
@@ 3.1 Merging induces an input-dependent target a single expert cannot realize
Define the router’s input-dependent mixing ratio $r(x):=\frac{g_{i}(x)}{g_{i}(x)+g_{j}(x)}\in[0,1]$ on the set where $g_{i}+g_{j}>0$. Substituting $g_{i}(x)$ and $g_{j}(x)$ in terms of $r(x)$, the original contribution of the pair $(i,j)$ can be written as
@@
After merging, the router must apply the summed gate, $g_{i}(x)+g_{j}(x)$, to a constant convex combination of the constituent experts which is independent of $x$. The core issue is that the merged model is forced to approximate the dynamic, input-dependent target expert with a static one. The following theorem quantifies this unavoidable approximation error.
Theorem 1 (Irreducible error of merging).
Let $\tilde{f}_{\alpha}(x)=\alpha f_{i}(x)+(1-\alpha)f_{j}(x)$ with a constant $\alpha\in[0,1]$ and define $\Delta_{ij}:=f_{i}(x)-f_{j}(x)$. The $L^{2}$ error of the merged pair is minimized when $\alpha$ is chosen to be the expected mixing ratio, $\alpha^{\star}:=\mathbb{E}[r(x)]$. Omitting the argument $(x)$ for brevity, this minimal error is
$\big\|\,\big(g_{i}{+}g_{j}\big)\!\big(rf_{i}{+}(1{-}r)f_{j}\big)\;-\;\big(g_{i}{+}g_{j}\big)\!\big(\alpha f_{i}{+}(1{-}\alpha)f_{j}\big)\big\|^{2}$ $=\underbrace{\mathbb{E}\!\left[(g_{i}{+}g_{j})^{2}\right]}_{\text{router scale}}\cdot\underbrace{\mathrm{Var}[r(x)]}_{\text{policy variability}}\cdot\underbrace{\|\Delta_{ij}\|^{2}}_{\text{expert gap}}.$ (5)
In particular, if the router’s policy is not constant ($\mathrm{Var}[r(x)]>0$) and the experts are not functionally identical ($\|\Delta_{ij}\|>0$), then every constant-$\alpha$ merge incurs strictly positive excess risk.
Proof.
The error term simplifies to $\big\|\big(g_{i}{+}g_{j}\big)\big(r-\alpha\big)\Delta_{ij}\big\|^{2}$. Assuming independence between the router policy and expert functions, this is proportional to $\mathbb{E}[(r-\alpha)^{2}]$. This is a standard least-squares problem minimized when $\alpha=\mathbb{E}[r]$, and the minimal value is $\mathrm{Var}[r]$. ∎
@@
Theorem 1 illustrates that merging with summed gates is fundamentally flawed whenever (i) the router has learned an input-dependent policy for mixing two experts ($\mathrm{Var}[r]>0$), and (ii) the experts are themselves distinct ($\|\Delta_{ij}\|>0$). Any fixed $\alpha$ cannot overcome the irreducible error bound established in equation 5.
@@ 3.2 Pruning preserves independent control
Pruning removes one function but importantly does not tie the remaining gates. The router still modulates each surviving expert independently. In contrast, merging removes a degree of freedom in the policy by replacing individual experts with their mergers. For a direct comparison under no fine-tuning, pruning expert $j$ and reallocating its gate-value to expert $i$ produces the error
$\big\|(g_{i}(x)f_{i}(x)+g_{j}(x)f_{j}(x))-(g_{i}(x){+}g_{j}(x))f_{i}(x)\big\|^{2}\;=\;\mathbb{E}\!\left[g_{j}(x)^{2}\,\|\Delta_{ij}(x)\|_{2}^{2}\right].$ (6)
Unlike equation 5, equation 6 does not penalize policy variability, the router still controls surviving experts independently. Whenever the router substantially mixes $i$ and $j$ (large $\mathrm{Var}[r]$) while the pruned expert $j$ has a small average gate-value ($\mathbb{E}[g_{j}^{2}]$), pruning admits a strictly smaller error than merging.
@@
Theorem 1 establishes that summed gate merging incurs an irreducible error directly proportional to the router’s policy variability ($\mathrm{Var}[r(x)]$). In contrast, the error from pruning a low-usage expert (Eq. 6) is proportional to its gate-value ($\mathbb{E}[g_{j}^{2}]$) and is insensitive to policy variability. Therefore, when the router actively mixes between two distinct experts, merging is fundamentally disadvantaged.
@@
(i) The constant-mixture model $\tilde{f}_{\alpha}$ is mathematically related to the frequency weighted parameter averaging merge used in practice. (ii) Even if $\tilde{f}$ was dependent on $x$, the router after merging cannot independently modulate the two latent functions, so the original policy is invalidated. (iii) With top-k routers, the specific irreducible error from policy variability ($\mathrm{Var}[r(x)]$) is generated exclusively on the support where both experts are selected. Outside that support, this component vanishes, leaving only a static error term that depends on the functional expert gap. (iv) See Appendix˜A for an extension of the above analysis to hierarchical clustering.

### v2

@@ Abstract
Sparsely-activated Mixture-of-Experts (SMoE) models offer efficient pre-training and low latency but their large parameter counts create significant memory overhead, motivating research into expert compression. Contrary to recent findings favouring expert merging on discriminative benchmarks, we find that expert pruning is a superior strategy for generative tasks. We demonstrate that existing merging techniques introduce an irreducible error due to the loss of fine-grained routing control over experts. Leveraging this insight, we propose Router-weighted Expert Activation Pruning (REAP), a novel pruning criterion that considers both router gate-values and expert activation norms to minimize the reconstruction error bound. Across a diverse set of SMoE models ranging from 20B to 1T parameters, REAP consistently outperforms merging and other pruning methods on generative benchmarks, especially at 50% compression. Notably, our method achieves near-lossless compression on code generation tasks with Qwen3-Coder-480B and Kimi-K2, even after pruning 50% of experts.
@@ 1 Introduction
Interest in the Sparsely-activated Mixture-of-Experts (SMoE) architecture for Large Language Models (LLMs) surged following the release of DeepSeek-V3 (DeepSeek-AI et al., 2024) and other high-quality open-weight SMoE LLMs (Jiang et al., 2024; Meta AI Team, 2025; Yang et al., 2025a; Zeng et al., 2025; Baidu, 2025; Kimi Team et al., 2025). Compared to dense models, SMoEs offer lower latency and more efficient pre-training (Fedus et al., 2022). However, SMoEs require more parameters than dense models to achieve similar accuracy, resulting in significant memory overhead. Further, expert usage imbalance during inference causes poor accelerator utilization, leading to increased latency or compromises such as dropped tokens (Balmau et al., 2025). Expert usage imbalance also represents an opportunity, motivating prior work which investigates whether experts can be compressed without negatively impairing accuracy (Li et al., 2023; Lu et al., 2024). By eliminating or compressing redundant experts, memory overhead is reduced. A more uniform distribution of expert usage would also improve hardware utilization. Expert compression is particularly valuable for use cases which feature small batch sizes such as local deployments and academic research.
@@
  • We demonstrate that existing expert merging techniques introduce irreducible error due to the loss of the router’s independent, input-dependent modulation of the expert outputs. In high-granularity SMoEs, the loss of fine-grained routing control results in functional subspace collapse;
  • Empirically, we find that expert merging distorts the functional manifold topology due to the introduction of novel functionality. Conversely, as a coordinate subspace operation, pruning preserves the topology;
  • We introduce Router-weighted Expert Activation Pruning (REAP), a novel expert pruning saliency criterion. By considering both router gate-values and expert activation norms, REAP explicitly minimizes the upper bound of the reconstruction error derived in our analysis by pruning experts which contribute minimally to the layer output;
  • Across diverse SMoE architectures ranging from 20B to 1T parameters and a suite of generative evaluations, we demonstrate the significant and consistent advantage of REAP over existing expert pruning and merging approaches, particularly at 50% compression. Notably, our method achieves near-lossless compression ($\Delta_{acc}\leq 2\%$) on code generation tasks after pruning 50% of experts from Qwen3-Coder-480B and Kimi-K2;
@@ 2 Related Work
A Mixture-of-Experts (MoE) layer is comprised of multiple, specialized feed-forward subnetworks known as experts and a router which produces gate-values (i.e., gates) to dynamically modulate the output of the experts based on the input. The architecture was revived in the deep learning era by the introduction of the SMoE by Shazeer et al. (2017). SMoEs layers only select a subset of experts to use for each input, enabling massive scaling of model parameters without a commensurate increase in computational cost (Lepikhin et al., 2021; Fedus et al., 2022). In transformer-based LLMs, SMoE layers are integrated by replacing the traditional feed-forward layers. Further innovations such as auxiliary-loss-free load balancing (DeepSeek-AI et al., 2024), shared experts, and fine-grained experts (Dai et al., 2024) have propelled SMoE architectures to become the de facto standard for LLMs in recent months.
@@
Although SMoE layers effectively decouple total model parameters from inference costs, the memory overhead of storing large SMoEs restricts their deployment in resource-constrained environments, motivating research in expert pruning to reduce total number of parameters. Early efforts demonstrated that progressively pruning experts based on router weights during fine-tuning until a single expert remained could preserve model quality in task-specific settings (Chen et al., 2022). Koishekenov et al. (2023) found expert pruning to be effective without further fine-tuning despite aggressively pruning up to 80% of experts.  Muzio et al. (2024) found that global pruning using gate-values as a saliency criterion was more effective than uniform, layer-wise frequency-based pruning. Other sophisticated pruning criteria have been proposed:  Lu et al. (2024) introduced an exhaustive search strategy which prunes experts that minimize the reconstruction loss between the original and pruned layer outputs;  Liu et al. (2024a) used a gradient-free evolutionary algorithm to prune experts. Both of these works demonstrated significant improvements over naive frequency-based pruning. A comprehensive evaluation of 16 diverse pruning criteria was conducted by Jaiswal et al. (2025). Expert Activation Norm (EAN) was empirically found to be the highest performing criterion and the benefits of iterative pruning were presented.
@@
While the above-noted works prove that expert compression is feasible via pruning, an alternative compression technique is to merge experts. Generally, merging requires both a clustering algorithm and a merging technique. Li et al. (2023) introduced Merge Sparse Mixture of Experts (M-SMoE) which first initializes expert cluster centres by identifying the dominant experts with the highest usage frequency globally across all layers. The remaining non-dominant experts are clustered based on the cosine similarity of router logits. Finally, expert weights are aligned via permutation with the weight matching algorithm (Ainsworth et al., 2023) and merged using frequency-weighted parameter averaging.  Li et al. (2023) found that their technique outperformed Chen et al.’s (2022) pruning method on MC benchmarks. Chen et al. (2025) proposed Hierarchical Clustering for Sparsely activated Mixture of Experts (HC-SMoE). HC-SMoE clusters experts based on the euclidean similarity of their representative vectors — the average activation of each expert measured on every token in a calibration dataset — using hierarchical agglomerative clustering. Similar to M-SMoE, HC-SMoE uses frequency-weighted parameter averaging to merge clusters into a single merged expert. Without any fine-tuning, Chen et al. (2025) found that their technique outperformed expert pruning based on router logits (He et al., 2025a), frequency, and Lu et al.’s (2024) method when benchmarked on a suite of MC question answering tasks.
@@
In addition to pruning and merging, experts may be compressed through quantization (Huang et al., 2025; Li et al., 2025; Duanmu et al., 2025), low-rank decomposition (Yang et al., 2024a; Gu et al., 2025; He et al., 2025b), weight sparsity (He et al., 2025a), or a combination of any of the above techniques (Liu et al., 2025a). These other approaches are orthogonal to expert pruning and merging; however, note that expert merging necessitates re-quantization for block quantization formats that share common scaling coefficients across a group of weights whereas pruning does not.
@@
3 Motivation
@@
To motivate our proposed expert pruning method, we derive the expected errors of both expert merging and pruning. Consider a SMoE layer with $K$ experts $f_{1},\dots,f_{K}$, each a function $f_{k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Let $\mathcal{T}(x)$ denote the set of indices corresponding to the top-$k$ router scores. The router produces a sparse gating vector $\mathbf{g}(x)\in\mathbb{R}^{K}_{\geq 0}$ where $g_{k}(x)>0$ if $k\in\mathcal{T}(x)$ and $g_{k}(x)=0$ otherwise. We assume the active gates are normalized such that $\sum_{k\in\mathcal{T}(x)}g_{k}(x)=1$, an operation commonly included in SMoE architectures. The output of the layer is
$h(x):=\sum_{k\in\mathcal{T}(x)}g_{k}(x)\,f_{k}(x).$ (1)
@@
[…]
$\bar{h}(x):=\sum_{k\neq j}\frac{g_{k}(x)}{1-g_{j}(x)}f_{k}(x),$ (2)
$\tilde{h}(x):=\big(g_{i}(x)+g_{j}(x)\big)\tilde{f}(x)+\sum_{k\neq i,j}g_{k}(x)f_{k}(x).$ (3)
@@ 3.1 Merging induces an input-dependent target a single expert cannot realize
Define the router’s input-dependent mixing ratio $r(x):=\frac{g_{i}(x)}{g_{i}(x)+g_{j}(x)}\in[0,1]$ locally on the set where $g_{i}+g_{j}>0$. Substituting $g_{i}(x)$ and $g_{j}(x)$ in terms of $r(x)$, the original contribution of the pair $(i,j)$ can be written as
@@
After merging, the router must apply the summed gate, $g_{i}(x)+g_{j}(x)$, to a constant convex combination of the constituent experts which is independent of $x$. The core issue is that the merged model is forced to approximate the dynamic, input-dependent target expert with a static one. The following quantifies this unavoidable approximation error.
Irreducible error of merging.
Let $\tilde{f}(x)=\alpha f_{i}(x)+(1-\alpha)f_{j}(x)$ with a constant $\alpha\in[0,1]$ and define $\Delta_{ij}:=f_{i}(x)-f_{j}(x)$. This definition of $\tilde{f}$ assumes that the experts are linear functions of $x$ which is generally not the case; however, this simplified model approximates the behaviour of frequency-weighted parameter averaging used by expert merging techniques in practice. $\mathcal{E}_{merge}$ is minimized when $\alpha$ is chosen to be the expected mixing ratio, $\alpha^{\star}:=\mathbb{E}[r(x)]$. Omitting the argument $(x)$ for brevity, this minimal error is
$\big\|\,\big(g_{i}{+}g_{j}\big)\!\big(rf_{i}{+}(1{-}r)f_{j}\big)\;-\;\big(g_{i}{+}g_{j}\big)\!\big(\alpha^{\star}f_{i}{+}(1{-}\alpha^{\star})f_{j}\big)\big\|^{2}$ $=\mathbb{E}_{x}\!\big[\underbrace{(g_{i}{+}g_{j})^{2}}_{\text{router scale}}\cdot\underbrace{(r-\alpha^{\star})^{2}}_{\text{policy variability}}\cdot\underbrace{\|\Delta_{ij}\|^{2}}_{\text{expert gap}}\big].$ (5)
In particular, if the router’s policy is not constant ($\mathrm{Var}[r(x)]>0$) and the experts are not functionally identical ($\|\Delta_{ij}\|>0$), then every constant-$\alpha$ merge incurs positive error. Let $G_{ij}:=\mathbb{E}_{x}[\lVert\Delta_{ij}(x)\rVert_{2}^{2}]$. Under a simplifying assumption that the router scale, policy variability, and $G_{ij}$ are weakly correlated across inputs, the error term may be decomposed to:
$\mathbb{E}_{x}[(g_{i}(x)+g_{j}(x))^{2}\,(r(x)-\alpha^{\star})^{2}\,\lVert\Delta_{ij}(x)\rVert_{2}^{2}]\approx\mathbb{E}_{x}[(g_{i}(x)+g_{j}(x))^{2}]\cdot\mathrm{Var}[r(x)]\cdot G_{ij}$ (6)
@@
This is a standard least-squares problem minimized when $\alpha=\mathbb{E}[r]$, and the minimal value is $\mathrm{Var}[r]$. Based on the assumptions noted above, we conclude that merging with summed gates is fundamentally flawed whenever: (i) the router has learned an input-dependent policy for mixing two experts ($\mathrm{Var}[r]>0$); and (ii) the experts are themselves distinct ($\|\Delta_{ij}\|>0$). Any fixed $\alpha$ cannot overcome the irreducible error bound established in Equation˜6.
@@ 3.2 Pruning preserves independent control
Pruning removes one function but importantly does not tie the remaining gates. The router still modulates each surviving expert independently. In contrast, merging removes a degree of freedom in the policy by replacing individual experts with their mergers. For a direct comparison under no fine-tuning, we consider the error of pruning expert $j$ where $j\in\mathcal{T}(x)$. After pruning, the router promotes previously inactive expert $i$ with the new gate-value of $g_{i}^{\prime}(x)\neq 0$, producing the error
$\mathcal{E}_{prune}=\mathbb{E}_{x|j\in\mathcal{T}(x)}\big[\big\|\underbrace{g_{j}(x)f_{j}(x)-g_{i}^{\prime}(x)f_{i}(x)}_{\text{substitution error}}-\underbrace{\frac{g_{j}(x)-g_{i}^{\prime}(x)}{1-g_{j}(x)+g_{i}^{\prime}(x)}\sum_{k\neq i,j}g_{k}(x)f_{k}(x)}_{\text{renormalization error}}\big\|^{2}_{2}\big]$ (7)
Substitution error is the dominant term in the above expression as the renormalization error coefficient simply scales the magnitude of the surviving expert outputs without changing their direction. In contrast, the substitution error includes the output of the promoted expert which may introduce significant error. With top-$k$ routing $g_{i}^{\prime}\leq g_{j}$ and the maximum substitution error occurs when $g_{i}^{\prime}\approx g_{j}$ with a magnitude upper bounded by
[…]
@@
While neither method is clearly superior for all distributions, our simplified analysis above isolates specific sources of error. Merging with summed gates couples the experts, incurring error whenever either expert is active, unless the experts are functionally identical ($\Delta_{ij}\approx 0$). The router loses the ability to independently modulate the merged experts in an input-dependent manner. Equation˜6 establishes that summed gate merging incurs an irreducible error directly proportional to the router’s policy variability ($\mathrm{Var}[r(x)]$).
In contrast, pruning only incurs errors when the pruned expert is in the top-$k$ set, $j\in\mathcal{T}(x)$. Unlike Equation˜5, Equation˜8 does not penalize policy variability; the router still controls surviving experts independently. The substitution error from pruning (Eq. 7) is proportional to its gate-value ($g_{j}$) and is insensitive to policy variability. Highly-granular SMoEs with many experts per layer use highly variable routing policies (high $\mathrm{Var}[r(x)]$) to combine many small contributions (small $g_{j}(x)$). In this setting, we expect merging with summed gates to be fundamentally disadvantaged.
@@
(i) The constant-mixture model $\tilde{f}$ is mathematically related to the frequency weighted parameter averaging merge used in practice. (ii) Even if $\tilde{f}$ was dependent on $x$, the router after merging cannot independently modulate the two latent functions, so the original policy is invalidated. (iii) With top-$k$ routers, the specific irreducible error from policy variability ($\mathrm{Var}[r(x)]$) is generated exclusively on the support where both experts are selected. Outside that support, this component vanishes, leaving only a static error term that depends on the functional expert gap. (iv) See Appendix˜A for an extension of the above analysis to hierarchical clustering.